# Affective Computing in the Era of Large Language Models: A Survey from the NLP Perspective

### v1

@@ 5.2.3 CoT for Affective Generation
In AG, information and knowledge are crucial for producing high-quality conversations. To obtain more valuable information, CoT is also introduced to assist affective generation. User states, serving as underlying linguistic cues, can reveal hidden needs, guiding the model to generate more personalized and engaging responses. The Cue-CoT framework [61] breaks down the reasoning process into several consecutive steps, progressively generating user state information such as personality, psychological state, and emotions based on the context, fully perceiving the linguistic cues within the context. The final step is to generate responses based on the previous reasoning output. The EDIT framework [175] also enhances dialogue response generation by detecting implicit user intentions. First, EDIT generates open-ended questions related to the context to capture potential user intentions. Then, it searches for answers in domain-specific knowledge bases to acquire additional knowledge. Finally, it enhances response generation by explicitly integrating this additional knowledge.
Since Affective Generation requires models to handle emotion-related issues in dialogues more proactively, planning strategic information becomes essential when utilizing LLMs to obtain information. The ProCoT [176] framework guides LLMs to generate descriptive information about intermediate steps in reasoning and planning, analyzing the following action towards the dialogue goal through dynamic reasoning and planning. Responses are generated based on this planning information, endowing the dialogue system with proactive dialogue capabilities. The ECoT framework [177] utilizes LLM to sequentially execute five steps: understanding context, recognizing others’ emotions, recognizing self-emotions, managing self-emotions, and influencing others’ emotions, aligning with human emotional intelligence to generate more emotionally rich responses. Additionally, to mimic the processes of recognizing, understanding, and regulating emotions, ESCoT [178] generates emotional stimuli and formulates strategies for emotion regulation to facilitate response generation.
Additionally, in the metaphor generation task, GROUNDS naturally serve as the CoT connecting TENOR and VEHICLE [179]. Therefore, researchers use the labels in annotated datasets as examples to guide the model in generating logical CoT content as GROUNDS. Ultimately, the model leverages these GROUNDS to generate coherent and comprehensive VEHICLES, thereby improving the quality of the generated metaphorical expressions.
@@ 5.2.4 Agent for Affective Generation
As a new technology based on LLMs, Agents are gradually being applied in AG. The Agent4SC system [55] generates user information, enhances memory capabilities, and formulates response strategies through the cooperation cooperation of agents, realistically simulating human behavior in using stickers, thereby enhancing multimodal empathetic dialogue. The COOPER dialogue framework [180] coordinates multiple specialized agents, each focusing on specific aspects of dialogue goals, to generate emotional support and persuasive dialogue responses.
Additionally, in the multimodal domain, researchers study multimodal Conversational Health Agents (CHAs) [181] and design five types of agents as components to interact with users to generate audio responses based on user emotions. These studies indicate that Agent has excellent potential in AG, effectively improving dialogue’snaturalness and empathetic capability.
@@ 6 Benchmark & Evaluation
This section compiles various benchmarks, evaluation methods, and metrics for Affective Computing, encompassing both Affective Understanding and Affective Generation. Table II comprehensively summarizes these benchmarks, detailing the evaluated tasks, datasets, metrics, and models.
TABLE II: Various benchmarks for Affective Computing. The AE means Automatic Evaluation, the HE means Human Evaluation.
@@
Text Annotation [96] Sentiment Analysis Sports tweets and Amazon reviews AE (Accuracy, F1, …) GPT-3.5-turbo
SOUL [114] Sentiment Analysis Yelp [182], IMDb [183] AE (F1, BLEU, …), HE Flan-T5, GPT-3.5-turbo
GPT-4V with Emotion[184] Sentiment Analysis MVSA-Single, DFEW, etc. 21 datasets AE (Accuracy, WAR, WAF) GPT-4V
MERBench [185] Multimodal Sentiment Analysis MER2023, etc. 7 datasets AE (WAF, MSE) PLMs, LLMs
MC-EIU [84] Emotion and Intent Joint Understanding in Multimodal Conversation MC-EIU test AE (WAF) EI2, other PLMs
LLMs2ESC [174] Emotional Support Conversation ESConv AE (Proficiency, Preference bias, ..) , HE GPT series, other LLMs
Cue-CoT [61] Emotional Support Conversation ED, etc. 6 datasets AE (Avg BLEU and F1), HE ChatGLM, other LLMs
ECoT [177] Emotional Support Conversation IEMCOP, COCO, et 7 datasets AE (EGS) GPT series, other LLMs
FEEL [186] Emotional Support Conversation ESCEval AE (Spear., Kend.), HE : GPT-3.5-turbo, ERNIE-4.0, GLM-4
MEDIC [187] Empathetic Response Generation MEDIC test AE (EE, ER, and CR) TFN, other PLMs
EmotionBench [188] Empathetic Response Generation EmotionBench test AE, HE GPT series, other LLMs
ChatGPT2ERG [189] Empathetic Response Generation EMPATHETICDIALOGUES AE (Distinct-n, BERTscore, BLEU-n), HE GPT-3.5-turbo, other PLMs
STICKERCONV [55] Multimodal Empathetic Response Generation STICKERCONV AE(MMr, BLEU, …), HE PEGS, ChatGLM3, Vicuna
SECEU [190] Emotional Intelligence Evaluation SECEU test AE (SECEU score, EQ score, …) GPT series, and other LLMs
EQ-Bench [113] Emotional Intelligence Evaluation EQ-Bench test AE (EQ-Bench scores) GPT series, other LLMs
EmoBench [191] Emotional Intelligence Evaluation EmoBench test AE (EmoBench scores) GPT series, other LLMs
EIBENCH [145] Emotional Intelligence, General Intelligence MMLU, etc. 88 datasets AE (Accuracy) Flan-T5, LLaMA-2
ChatGPT2AC [99] Affective Computing res14, et 13 datasets AE (Accuracy, UAR) gpt-4
MM-INSTRUCTEVAL [192] Multimodal Reasoning with Multimodal Contexts MVSA-Single etc. 16 datasets AE (Accuracy) GPT series, other LLMs
@@ 6.1 Affective Understanding
This section introduces the benchmarks in the field of Affective Understanding for different datasets related to SA.
Text Annotation [96] evaluates GPT-3.5-turbo as a tool for sentiment analysis labeling on two distinct datasets: sports tweets and Amazon reviews. GPT-3.5-turbo shows promise as a tool for text labeling for various sentiment analysis events and tasks.
Different from merely predicting the polarity of sentiment, a novel task called Sentiment and Opinion Understanding of Language (SOUL) is introduced, including two subtasks: Review Comprehension (RC) and Justification Generation (JG) [114]. RC seeks to validate statements that focus on subjective information based on a review text, while JG requires models to explain for their sentiment predictions.
GPT-4V with Emotion [184] presents the quantitative evaluation results of GPT-4V on 21 benchmark datasets covering six tasks: visual sentiment analysis, tweet sentiment analysis, micro-expression recognition, facial emotion recognition, dynamic facial emotion recognition, and multimodal emotion recognition. This paper collectively refers to these tasks as “Generalized Emotion Recognition (GER)”.
MERBench [185] is a benchmark for Multimodal Sentiment Analysis, covering of seven datasets: IEMOCAP [193], MELD [131], CMU-MOSI [194], CMU-MOSEI [195], CH-SIMS [196], CH-SIMS V2 [197]. MER2023, consisting of 3 datasets: MER-MULTI, MER-NOISE, MER-SEMI, is proposed as a Chinese benchmark. MERBench uses Automatic Evaluation, which includes the weighted average F-score (WAF) and the mean squared error (MSE).
MC-EIU dataset [84] is the first comprehensive and rich multimodal conversational emotion and intent joint understanding dataset, consisting of 4,970 conversational video clips from TV series in Chinese and English. It uses Automatic Evaluation, the Weighted Average F-score (WAF), as the metric. Emotion and Intent Joint Understanding in Multimodal Conversation (MC-EIU) aims to decode the semantic information manifested in a multimodal conversational history while simultaneously inferring the emotions and intents for the current utterance.

### v2

@@ 5.2.3 CoT for Affective Generation
In AG, information and knowledge are crucial for producing high-quality responses. To obtain more valuable information, researchers introduce CoT to support empathetic generation. CFEG Chen et al. (2024b) decomposes emotion and reasoning into multiple steps, strengthening the listener’s role awareness and significantly improving the empathy of LLMs. User utterances serve as latent linguistic cues that reveal hidden needs and guide the model to generate more personalized responses. Cue-CoT Wang et al. (2023a) breaks down reasoning into sequential steps to infer user states (e.g., personality, psychological status, emotions) and finally generates responses based on the reasoning outputs. EDIT Wu et al. (2023) enhances response generation by detecting implicit user intentions. It first asks open-ended questions to capture potential intentions, then retrieves knowledge base answers, and finally integrates the additional knowledge into generation. Aptness Hu et al. (2024) adopts a similar strategy by constructing an empathetic response database via multi-step CoT and using retrieval during generation to enhance empathetic capability.
Since AG requires LLMs to address emotion-related issues more proactively, planning information plays an important role in the process of obtaining and using information with LLMs. ProCoT Deng et al. (2023b) guides LLMs to produce intermediate steps of reasoning and planning and analyzes through dynamic reasoning how to achieve dialogue goals, which enhances proactive dialogue ability. ECoT Li et al. (2024g) proposes five consecutive steps to simulate human emotional intelligence and generate responses with richer emotional expression. ESCoT Zhang et al. (2024e) promotes response generation by producing emotional stimuli and designing regulation strategies. Further research transforms step-by-step cognitive commentary prompts into a hierarchical concept Li et al. (2024d) and proposes a three-level emotional integration framework.
Additionally, in metaphor generation, GROUNDS naturally serve as the reasoning chain that connects TENOR and VEHICLE. Researchers guide the model to generate logically consistent GROUNDS Shao et al. (2024a), which in turn produce coherent and rich VEHICLE expressions, thereby improving the quality of metaphor generation.
@@ 5.2.4 Agent for Affective Generation
As a new technology based on LLMs, agent is gradually applied to AG. The Agent4SC Zhang et al. (2024f) generates user information, enhances memory, and formulates response strategies through collaboration among LLMs, realistically simulating human sticker usage to improve multimodal empathetic dialogue. The COOPER dialogue framework Cheng et al. (2024a) coordinates multiple specialized agents focusing on specific dialogue goals to generate emotionally supportive and persuasive responses. Researchers propose a multi-turn dialogue framework based on psychologist agents Wu et al. (2024c), in which debaters from different psychological schools generate candidate replies, and an unbiased decision-maker selects the final response, thus addressing the challenge of integrating multiple psychological schools in LLMs. The PerceptiveAgent Yan et al. (2024) simulates different speaking styles according to prompts to build multimodal dialogue systems. The conversational agent with acoustic emotion perception (CA) Hu et al. (2022b) performs speech emotion recognition and incorporates empathetic feedback and interjections to give responses an emotional style. In addition, introducing self-emotion prompts is also shown to help Agents demonstrate more human-like dialogue strategies Zhang et al. (2024c).
Additionally, in the multimodal domain, researchers study multimodal Conversational Health Agents (CHAs) Abbasian et al. (2024) and design five types of agents as components to interact with users to generate audio responses based on user emotions.
@@ 5.3 Challenge
6 Reinforcement Learning
Reinforcement learning (RL) is increasingly used to align large language models (LLMs) with affect-related objectives beyond supervised fine-tuning. We organize RL for AC into three families that differ by the source of rewards yet share the goal of optimizing behavior at token-, utterance-, and dialogue-levels across AU/AG (see Figure 5): (i) Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022); Rafailov et al. (2023b), where a reward model trained on human comparisons guides policy optimization—often with Proximal Policy Optimization (PPO)—and preference-only variants such as Direct Preference Optimization (DPO) reduce reliance on online RL Ouyang et al. (2022); Schulman et al. (2017); Rafailov et al. (2023b); (ii) Reinforcement Learning with Verifiable/Programmatic Rewards (RLVR) Lambert et al. (2024), which replaces preference labels with rule- or checker-based signals and has inspired algorithms like Group Relative Policy Optimization (GRPO) Shao et al. (2024b) for stable group-wise ranking Lambert et al. (2024); Shao et al. (2024b); and (iii) Reinforcement Learning from AI Feedback (RLAIF), which leverages capable AI evaluators to provide multi-aspect feedback that can be plugged into RL or preference-optimization pipelines. The following subsections detail representative methods and applications of these three families in affective computing.
Figure 5: The overview of reinforcement learning.
6.1 Reinforcement Learning from Human Feedback
A key application of RLHF is enhancing empathy in dialogue systems. Zhu et al. Zhu et al. (2023a) propose a three-stage framework for Empathetic Emotion Elicitation, where reinforcement learning leverages user reactions as rewards to improve emotional engagement across conversations. However, collecting human feedback is costly and difficult to scale. To mitigate this, Xu et al. Xu et al. (2025) introduce Reinforcement Learning from Targeted Human Feedback (RLTHF), which first uses a general-purpose LLM to label data, then directs limited human effort toward “hard” or mislabeled cases. This approach achieves full human-level alignment at only 6–7% of the annotation cost.
The quality of preference data remains critical. Liang et al. Liang et al. (2024) present AlignCap, a preference optimization method that aligns speech emotion captions with human preferences, reducing hallucinations and producing more faithful emotional descriptions. Complementarily, Chen Chen et al. (2025) release EmpathyAgent, the first benchmark for training and evaluating empathetic embodied agents, offering 10,000 multimodal samples with preference-labeled empathetic actions for RLHF training.
6.2 Reinforcement Learning with Verifiable Rewards
To bypass the human annotation bottleneck, RLVR leverages programmatic or model-based reward functions, enabling scalable and automated optimization. A representative work is RL-EMO by Zhang et al. Zhang et al. (2024a), the first to apply RL to multimodal emotion recognition in conversation (ERC). It frames ERC as a sequential decision-making problem, where correct or incorrect emotion predictions yield positive or negative rewards, respectively. Using a Deep Q-Network to model emotion flow, RL-EMO outperforms prior methods and demonstrates the effectiveness of policy optimization for capturing temporal affective dynamics. Zhang et al. Zhang & Jaitly (2025) further propose SAGE (State Augmented GEneration), a framework that explicitly models user and assistant emotional states as latent variables for dialogue generation. Reinforcement learning is applied at the state level, enabling strategic emotional control and producing more empathetic and context-aware responses. Extending into the multimodal domain, Zhao et al. Zhao et al. (2025a) present R1-Omni, which applies RLVR to omni-multimodal (video, audio, text) emotion recognition. With a simple verifiable reward—whether predictions match ground-truth labels—optimized via Group Relative Policy Optimization (GRPO), the framework improves accuracy, robustness, and interpretability in emotion reasoning.
Pushing the boundaries of affective understanding beyond static emotion recognition, the SAGE (Sentient Agent as a Judge) framework Zhang et al. (2025a) reconceptualizes comprehension as a dynamic, interactive process. Its core innovation is a "Sentient Agent" that simulates human-like emotional changes and inner thoughts during multi-turn conversations. By embedding evaluation within the agent’s evolving persona, goals, and emotional state, comprehension is reframed from classification to a stateful simulation of a user "feeling understood." The resulting emotion trajectories and interpretable inner thoughts provide a proxy for modeling and measuring the essence of being heard.
Most RL applications in affective computing focus on generation, steering LLMs toward empathetic, polite, or strategically effective outputs. RLHF naturally fits this domain by embedding human preferences into the training loop. For example, RLVER by Wang et al. Wang et al. (2025b) introduces the first end-to-end RL framework leveraging verifiable emotion rewards from simulated users. Built upon the SAGE simulator, it produces deterministic emotion scores as direct PPO rewards, elevating a 7B open-source model to near-proprietary performance on Sentient-Benchmark. A key innovation in RLVR is finer-grained reward design. Li et al. Li et al. (2024e) propose TOLE, which formulates token-level rewards via classifier probability shifts, enabling efficient controllable text generation. Cai Cai et al. (2024)’s EmpCRL combines empathy and diversity signals, while Li et al. Li et al. (2024c) integrate the "Cognitive Relevance Principle" into ORL, balancing helpfulness against user effort. Together these works show how automated, dense rewards lead to nuanced affective generation. Ma et al. Ma et al. (2025) propose EmpRL, aligning response empathy levels through a programmatic reward from a dedicated "empathy identifier," based on reaction, interpretation, and exploration. Chang et al. Chang et al. (2024) further demonstrate multi-objective reward design in emotional support systems, combining goal-level, flow, and valence rewards for stage-aware dialogue policies. RLVR also supports domain-specific agents. Mishra et al. Mishra et al. (2024) introduce ABLE, a disability support agent guided by a six-component reward model covering persona consistency, politeness, empathy, naturalness, and coherence. Priya et al. Priya et al. (2025) develop GENTEEL-NEGOTIATOR, which uses MoE-based RL with detailed reward functions for polite yet strategic negotiation. For online safety, Wang et al. Wang et al. (2024d) present F2RL, leveraging triplet-based factuality and multi-aspect faithfulness rewards to generate effective counterspeech.
Beyond dialogue, RLVR principles extend into strategic decision-making. Unnikrishnan et al. Unnikrishnan (2024) apply sentiment-aware RL to portfolio management, where sentiment scores from financial news served as programmatic rewards. The resulting trading agent achieved superior profitability, showing how effective signals can optimize complex, non-dialogue tasks.
6.3 Reinforcement Learning from AI Feedback
Within the RLVR paradigm, Reinforcement Learning from AI Feedback (RLAIF) replaces human annotators with capable AI models to provide reward signals. Yoshida et al. Yoshida et al. (2025) exemplify this approach by training an LLM-based reward model on 12 conversational metrics (e.g., empathy, consistency, enjoyability). This AI-generated, multi-faceted feedback is then used with DPO to optimize dialogue agents, showing that AI feedback can scalably capture holistic qualities beyond turn-level scoring.
RLVR also supports new agent architectures. Yuan Yuan et al. (2024) propose ReflectDiffu, integrating RL into diffusion models where a policy network selects empathetic intents conditioned on emotional contagion and mimicry. High-level planning approaches include Deng Deng et al. (2023c)’s Plug-and-Play Policy Planner and Rakib Rakib et al. (2025)’s DialogXpert, which employ LLM critics in self-play to train efficient policy planners for proactive tasks such as emotional support and negotiation. The paradigm further extends to autonomous self-improvement. Ye et al. Ye et al. (2025) propose a self-evolution framework where models refine their own outputs to generate synthetic preference pairs for DPO. Zhao et al. Zhao et al. (2025b) advance this idea with Chain-of-Strategy Optimization (CSO), which uses Monte Carlo Tree Search to explore conversational strategies. Candidate responses are scored by an auxiliary reward model, yielding high-quality preference datasets for DPO fine-tuning, significantly improving strategic empathy. Zhang et al. Zhang et al. (2025b) introduce Decoupled ESC, featuring Inferential Preference Mining (IPM) that identifies psychological errors (e.g., lack of empathy, premature shifts) in model outputs. Erroneous responses are paired with ground-truth ones to build synthetic preference data. The resulting decoupled system, with separate strategy planning and response generation, yields more robust and effective emotional support agents. Collectively, these works illustrate the trajectory of RLAIF toward fully autonomous systems capable of continuously enhancing emotional intelligence.
6.4 Challenge
Despite the growing success of reinforcement learning in aligning LLMs with affective objectives, several key challenges remain:
  • While RLHF has significantly improved model alignment with human preferences, it remains highly resource-intensive, requiring large-scale, high-quality human preference data. This poses a major scalability bottleneck, especially for fine-grained affective tasks that require nuanced and subjective judgments.
  • Many current reward functions used in RLHF and RLVR are designed based on task-level or surface-level attributes (e.g., helpfulness, politeness), but they often lack grounding in psychological or cognitive theories of emotion. This can lead to misalignment between the reward signal and the underlying affective dynamics, limiting the depth of emotional understanding or expression the model can achieve.
  • Despite advances in RL for affective understanding (e.g., RL-EMO, SAGE), temporal affect modeling remains underexplored. Emotions evolve over time, yet most RL-based affective systems still optimize at the turn-level or token-level, failing to capture long-term emotional trajectories in conversation.
  • The lack of standardized benchmarks and evaluation metrics for affective RL makes it difficult to systematically compare approaches or track progress. Current evaluations often rely on subjective human scoring or limited automatic proxies, which may not fully reflect emotional richness, coherence, or alignment with human expectations.
@@ 7 Benchmark & Evaluation
This section compiles various benchmarks, evaluation methods, and metrics for Affective Computing, encompassing both Affective Understanding and Affective Generation. Table 2 comprehensively summarizes these benchmarks, detailing the evaluated tasks, datasets, metrics, and models.
Table 2: Various benchmarks for Affective Computing. The AE means Automatic Evaluation, the HE means Human Evaluation.
@@
SOUL Deng et al. (2023d) Sentiment Analysis Yelp Zhang et al. (2015), IMDb Maas et al. (2011) AE (F1, BLEU, …), HE Flan-T5, GPT-3.5-turbo
GPT-4V with Emotion Lian et al. (2023b) Sentiment Analysis MVSA-Single, DFEW, etc. 21 datasets AE (Accuracy, WAR, WAF) GPT-4V
MERBench Lian et al. (2024) Multimodal Sentiment Analysis MER2023, etc. 7 datasets AE (WAF, MSE) PLMs, LLMs
DFEW  Jiang et al. (2020) Videol Sentiment Analysis 1500 movies AE(UAR,WAR) EC-STFL
MC-EIU  Liu et al. (2024b) Emotion and Intent Joint Understanding in Multimodal Conversation MC-EIU test AE (WAF) EI2, other PLMs
RLVER Wang et al. (2025b) Empathetic Understanding Sentient Benchmark AE (Sentient Benchmark) GPT-4, GPT-4V, DeepSeek, Claude3.7
SAGE Zhang et al. (2025a) Empathetic Understanding 100-scenario Supportive-Dialogue benchmark AE(BLRI, Sentient emotion score) GPT-4o, Gemini2.5
Cue-CoT Wang et al. (2023a) Emotional Support Conversation ED, etc. 6 datasets AE (Avg BLEU and F1), HE ChatGLM, other LLMs
ESC-EvalZhao et al. (2024a) Emotion Support Conversations 2,801 Role Cards Benchmark AE (ESC-RANK)& HE (7 dimensions) GPT-4, Llama3, EmoLLM, etc.
EmotionBench tse Huang et al. (2023) Empathetic Response Generation EmotionBench test AE, HE GPT series, other LLMs
STICKERCONV Zhang et al. (2024f) Multimodal Empathetic Response Generation STICKERCONV AE(MMr, BLEU, …), HE PEGS, ChatGLM3, Vicuna
SECEU Wang et al. (2023d) Emotional Intelligence Evaluation SECEU test AE (SECEU score, EQ score, …) GPT series, and other LLMs
EQ-Bench Paech (2024) Emotional Intelligence Evaluation EQ-Bench test AE (EQ-Bench scores) GPT series, other LLMs
EmoBench Sabour et al. (2024) Emotional Intelligence Evaluation EmoBench test AE (EmoBench scores) GPT series, other LLMs
EIBENCH Zhao et al. (2024b) Emotional Intelligence, General Intelligence MMLU, etc. 88 datasets AE (Accuracy) Flan-T5, LLaMA-2
EmoVIT Xie et al. (2024) Image Sentiment Explaination Generation EmoVIT test AE (Accuracy) EmoVIT, other MLLMs
EMER Lian et al. (2023a) Explainable Multimodal Emotion Recognition MER2023 AE(Accuracy, BLEU, ROUGE) EMER model, other MLLMs
MER-UniBench Lian et al. (2025) Multimodal Emotion Recognition and Reasoning MER-Caption AE(UAR,WAR), HE AffectGPT, other MLLMs
ChatGPT2AC Amin et al. (2024) Affective Computing res14, et 13 datasets AE(Accuracy, UAR) GPT-4
MM-INSTRUCTEVAL Yang et al. (2024b) Multimodal Reasoning with Multimodal Contexts MVSA-Single etc. 16 datasets AE (Accuracy) GPT series, other LLMs
@@ 7.1 Affective Understanding
This section introduces the benchmarks and representative datasets in the field of Affective Understanding for different sentiment analysis-related tasks.
SOUL Deng et al. (2023d) introduces two subtasks: Review Comprehension (RC) to validate subjective statements from reviews, and Justification Generation (JG) to explain sentiment predictions.
GPT-4V with Emotion Lian et al. (2023b) evaluates GPT-4V’s capabilities in Generalized Emotion Recognition (GER), providing quantitative results across 21 benchmark datasets spanning six visual and multimodal emotion recognition tasks.
MERBench Lian et al. (2024) is a benchmark for Multimodal Sentiment Analysis, covering seven prominent datasets. It uses Automatic Evaluation with metrics like weighted average F-score (WAF) and mean squared error (MSE).
DFEW Jiang et al. (2020) is a large-scale dataset for dynamic facial expression recognition, containing over 16,000 video clips from movies with real-world challenges like occlusions and variable poses. Evaluation metrics include UAR and WAR.
MC-EIU Liu et al. (2024b) is a multimodal conversational dataset with nearly 5,000 video clips from bilingual TV series. It targets joint emotion and intent understanding in conversations, evaluated using weighted average F-score (WAF).
RLVER Wang et al. (2025b) is a reinforcement learning framework for enhancing empathetic reasoning in LLMs. It uses the Sentient Benchmark to evaluate a model’s ability to improve a simulated user’s emotional state over time through verifiable emotion rewards.
SAGE Zhang et al. (2025a) is an evaluation framework for higher-order social cognition in LLMs. It uses a simulated Sentient Agent to track emotional changes during dialogues, evaluating a model’s ability to foster positive emotional shifts via its primary metric, the Sentient emotion score.